# algorithmic

### v1

@@ 6 Conclusion
In this paper, we present ARROW-Diff, a novel approach for large-scale graph generation based on random walk diffusion. ARROW-Diff generates graphs by integrating two components: (1) An order agnostic autoregressive diffusion model on the level of random walks that learns the generative process of random walks of an input graph, and (2) a GNN component that learns to filter out unlikely edges from the generated graph. Due to the random-walk based diffusion, ARROW-Diff efficiently scales to large graphs, significantly reducing generation time compared to baselines. It generates graphs of high quality, reflected in the high performance on many graph statistics. One limitation of ARROW-Diff is that it generates graphs with the same number of nodes as the original graph, due to the framework of the OA-ARDM. Potential future work could focus on a more flexible implementation of ARROW-Diff including an adaptation for learning on multiple graphs.
Appendix A Standard Deviations of All Evaluation Metrics
@@
ARROW- 9 0.002 320 0.02 0.007 0.000 1.5% 0.0
Diff
ARROW- 13 0.004 326 0.01 0.006 0.000 1.5% 0.0
Diff (w/o features)
@@
ARROW- 15 0.003 2,040 0.00 0.002 0.000 0.7% 0.7
Diff
ARROW- 35 0.002 4,714 0.01 0.003 0.000 1.3% 1.3
Diff (w/o features)
@@
ARROW- 5 0.007 69 0.03 0.008 0.000 1.4% 0.0
Diff
ARROW- 8 0.008 68 0.02 0.008 0.000 1.1% 0.2
Diff (w/o features)
@@
ARROW- 26 0.002 2,202 0.01 0.002 0.000 0.8% 0.7
Diff
ARROW- 27 0.003 2,797 0.01 0.002 0.000 0.6% 0.0
Diff (w/o features)
@@
ARROW- 11 0.003 1,454 0.00 0.001 0.000 0.8% 1.1
Diff
ARROW- 15 0.004 2,427 0.01 0.001 0.000 1.0% 0.0
@@
Appendix B Visualization of Community Structures Captured by ARROW-Diff
(a) Graphs from the test split.
(b) Graphs graphs from GraphRNN.
(c) Generated graphs from EDGE.
(d) Generated graphs from ARROW-Diff.
Figure 5: Visualization of six generated graphs from GraphRNN, EDGE, and ARROW-Diff trained on the Community-20 dataset, showing that ARROW-Diff is able to capture community structures in the ground-truth graph.
Appendix C Node Degree Distribution of Generated Graphs
Figure 6: Node degree distribution of the original CiteSeer graph and one generated graph from ARROW-Diff using the node features from CiteSeer.

### v2

@@ 7 Conclusion
In this paper, we present ARROW-Diff, a novel approach for large-scale graph generation based on random walk diffusion. ARROW-Diff generates graphs by integrating two components into an iterative procedure: (1) An order agnostic autoregressive diffusion model on the level of random walks that learns the generative process of random walks of an input graph, and (2) a GNN component that learns to filter out unlikely edges from the generated graph. Due to the random walk-based diffusion, ARROW-Diff efficiently scales to large graphs, significantly reducing the generation time compared to baselines. In our experiments, we show how ARROW-Diff generates graphs of high quality, reflected in the high performance on many graph statistics.
Acknowledgements
Tobias Bernecker and Ghalia Rehawi are supported by the Helmholtz Association under the joint research school “Munich School for Data Science - MUDS”. Francesco Paolo Casale was funded by the Free State of Bavaria’s Hightech Agenda through the Institute of AI for Health (AIH). Janine Knauer-Arloth’s contributions were supported by the Brain & Behavior Research Foundation (NARSAD Young Investigator Grant, #28063). Annalisa Marsico acknowledges support by the BMBF Cluster4Future program CNATM.
Appendix A Standard Deviations of the Results in Table 2
@@
BiGG 45 0.029 65 0.02 0.004 0.003 0.1% 1.8
ARROW-Diff 9 0.002 320 0.02 0.007 0.000 1.5% 0.0
ARROW-Diff 13 0.004 326 0.01 0.006 0.000 1.5% 0.0
(w/o features)
@@
BiGG 73 0.028 1,042 0.05 0.001 0.000 0.1% 127.9
ARROW-Diff 15 0.003 2,040 0.00 0.002 0.000 0.7% 0.7
ARROW-Diff 35 0.002 4,714 0.01 0.003 0.000 1.3% 1.3
(w/o features)
@@
BiGG 0 0.000 1 0.00 0.001 0.000 0.1% 1.1
ARROW-Diff 5 0.007 69 0.03 0.008 0.000 1.4% 0.0
ARROW-Diff 8 0.008 68 0.02 0.008 0.000 1.1% 0.2
(w/o features)
@@
BiGG 12 0.022 57 0.01 0.000 0.000 0.1% 11.7
ARROW-Diff 26 0.002 2,202 0.01 0.002 0.000 0.8% 0.7
ARROW-Diff 27 0.003 2,797 0.01 0.002 0.000 0.6% 0.0
(w/o features)
@@
BiGG 7 0.010 3 0.16 0.000 0.000 0.0% 32.5
ARROW-Diff 11 0.003 1,454 0.00 0.001 0.000 0.8% 1.1
ARROW-Diff 15 0.004 2,427 0.01 0.001 0.000 1.0% 0.0
(w/o features)
Appendix B Evaluation of the Quality of the Generated Graphs
In the following analysis, we validate the capability of ARROW-Diff in learning the topological structure of the original graph and in generating new graphs that have a similar data distribution as the original graph. To this end, we generated 100 graphs for each of CiteSeer, DBLP, and PubMed graphs. We used the graph embedding technique Graph2Vec (Narayanan et al., 2017) to map the generated graphs into embeddings of size 256. A principal component analysis (PCA) of the embeddings (Figure 5) revealed a clear separation between the different generated graphs of the different datasets highlighting the ability of ARROW-Diff in generating graphs with similar structural properties for the same dataset. The results also depicted that the embeddings of the generated graphs are close in the embedding space to their original counterparts, indicating that ARROW-Diff produces graphs that closely mimic the structural properties of the input graphs and underscoring the fidelity of the generated graphs. A logistic regression-based classifier trained on the generated graphs and tested on the real graphs reported a Classification Accuracy Score (CAS) (Ravuri & Vinyals, 2019) of 0.67.
Figure 5: Principal component analysis (PCA) of the Graph2Vec embeddings of the original CiteSeer, DBLP, and PubMed graphs and 100 generated graphs for each of the datasets, showing a clear separation between the generated graphs of the different datasets, and a proximity of the generated graphs to their real graph counterparts.
Appendix C Supplementary Experiments and Results
C.1 Stochastic Block Model Dataset
In the following experiment, we evaluate the ability of ARROW-Diff to generate graphs with a different structure to those shown in Section 4. Specifically, we create a synthetic dataset containing a Stochastic Block Model (SBM) (Holland et al., 1983) graph that does not show a preferential attachment property (power-law degree distribution). The constructed SBM graph includes three communities with 60, 100, and 200 nodes, and a total of 3,824 edges. Similar to our main experiments (Section 4), we split the edges into training, validation, and test edges, and use the same training split across all experiments.
C.2 Parameters for ARROW-Diff Training and Graph Generation
Due to the smaller size of the SBM graph, we set the random walk length to $D=8$ and run ARROW-Diff for $L=5$ iterations only. We also set the dimensionality of positional encodings (used by the GNN component) of the nodes to 64.
Table 5: The results of ARROW-Diff and six baseline models on the SBM dataset with 360 nodes and 3,824 edges.
Dataset Max. Assort- Triangle Avg. Global Edge Time
[…]
SBM 36 -0.016 3,623 0.134 0.019 - -
[…]
VGAE 122 -0.012 146,304 0.383 0.013 22.1% 0.0
Graphite 11 -0.134 23 0.048 0.082 0.3% 0.1
EDGE 34 0.017 1,581 0.072 0.012 5.7% 1.6
GraphRNN 14 -0.061 57 0.025 0.038 1.7% 0.2
BiGG 33 -0.029 2,204 0.113 0.019 65.9% 13.9
ARROW- 48 0.026 2,969 0.157 0.020 40.5% 0.5
@@ C.2 Parameters for ARROW-Diff Training and Graph Generation
C.3 Results
Table 5 presents the results as the mean across 10 generated graphs. For evaluation, we omit the power-law exponent metric, as it is exclusively applicable to scale-free graphs. The results demonstrate that ARROW-Diff achieves the best performance on the triangle count metric, consistent with the results on the citation graphs (Table 2). It also shows competitive performance compared to other baselines on the rest of the metrics, performing second best on two metrics while significantly reducing graph generation time. In Figure 6, we provide a visualization of the training split of the synthetic SBM graph as well as the generated graphs, showcasing the ability of ARROW-Diff in capturing the three community structures of the SBM graph.
Figure 6: Visualization of the SBM training graph as well as generated graphs from NetGAN, Graphite, EDGE, BiGG, and ARROW-Diff. We observe that ARROW-Diff is able to capture the community structures of the original graph.
Appendix D Node Degree Distribution of Generated Graphs
(a) CiteSeer dataset.
(b) SBM dataset.
Figure 7: Node degree distributions of the real CiteSeer and SBM graphs and one generated graph of ARROW-Diff for each of the two datasets. The original node features are used by ARROW-Diff for the CiteSeer dataset.